OpenStudy (fatemahzahrah):

sin 4x=1 ; I need all the solutions.... plz help

OpenStudy (faiqraees):

Range?

OpenStudy (fatemahzahrah):

yes

OpenStudy (faiqraees):

I am asking the range

OpenStudy (fatemahzahrah):

0<x<pi

OpenStudy (fatemahzahrah):

help me plz

OpenStudy (jpharper12345678cars):

sin4x=1/2 4x= π/6 , 5π/6 , 13π/6 , 17π/6 , 25π/6 , 29π/6 , 37π/6 , 41π/6 these are the points where sinx is equal to 1/2. the 4 in the original problem means you need to go around the until circle 4 times. so you would add 2π to each radian. to get x by itself you divide each of those answers by 4 so, example π/6 / 4 .... π/6 * 1/4 = π/24 so your answers in radians are π/24, 5π/24, 13π/24, 17π/24, 25π/24, 29π/24, 37π/24, and 41π/24 i hope that helps!!!!!!!!!! :) its confusing but i hope you understand it well!

OpenStudy (jpharper12345678cars):

did that help?

OpenStudy (faiqraees):

Hi do you still want help?

OpenStudy (fatemahzahrah):

it helps... thank u

OpenStudy (anonymous):

\[\sin4x=1=\sin(2n \pi+\frac{ \pi }{ 2 })\] \[4x=2n \pi+\frac{ \pi }{ 2 },x=\frac{ 4n+1 }{ 8 }\pi\] where n is an integer \[x=\frac{ \pi }{ 8 },\frac{ 5 \pi }{ 8 },~for~0<x<\pi\]